A 0.5 hp, 6-pole induction motor is excited by a 3 –phase, 60 Hz source. If the full-load speed is 

1140 rpm, calculate the slip.


1
Expert's answer
2021-09-15T18:36:29-0400

We proceed to calculate the speed of the rotor field with respect to the stator and then we can calculate the slip s:


"n_s=\frac{120f}{p}=\frac{120(60)}{6}=1200\,rpm
\\ s=\cfrac{n_s-n}{n_s}=\cfrac{1200-1140}{1140}
\\ \implies s=0.0526=5.26\, \%"


In conclusion, under these circumstances, the slip is s=0.0526=5.26 %.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.